OpenStudy (anonymous):

Factor completely. 3tx+6t+3hx+6h

OpenStudy (anonymous):

group the first and third

OpenStudy (anonymous):

group the second and fourth

OpenStudy (anonymous):

3tx+6t+3hx+6h=3t(x+2)+3h(x+2)=(x+2)(3t+3h)=(x+2)3(t+h)

OpenStudy (anonymous):

which is the answer?

OpenStudy (anonymous):

I think what I typed is correct. Perhaps, the last statement is the final answer.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

You're most welcome.